# 37N35, 65F55, 93A15, 93B15, 93C57

### v1

@@ 3 Data-driven second-order balanced truncation
where $\boldsymbol{\phi}(s)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\left(s\boldsymbol{M}+\boldsymbol{K}\right)$. To prove eqs. 22a and 22b, we use the resolvent identities from Lemma 1. Under the assumption that $h(\mathfrak{i}\vartheta_{k})\neq h(\mathfrak{i}\zeta_{j})$ for all $k$ and $j$, applying eq. 19a reveals that
@@
Replacing the intrusive quantities eq. 14 in Algorithm 1 with the data-based approximations eq. 21 results in a data-driven reformulation of $\mathsf{sopvBT}$ summarized in Algorithm 2, which we call second-order quadrature-based position-velocity balanced truncation ($\mathsf{soQuadpvBT}$). Outside of empirical knowledge used to design the damping coefficient functions $f(s)$ and $g(s)$, the method is entirely non-intrusive. As in [21], the quadrature-based approximations to the Gramians are never explicitly formed. Instead, they are only invoked implicitly to derive the data-based approximations eq. 21. While Algorithm 2 is formulated generally to include position and velocity outputs, as is the case with Theorem 1, both are not necessarily required; depending on the application of interest, only one of them is required. Lastly, Theorem 1 assumes that the sets of quadrature nodes used to implicitly approximate $\boldsymbol{P}_{\operatorname{p}}$ and $\boldsymbol{Q}_{\operatorname{v}}$ are distinct. In Theorem 2, we derive a special case of the formulae eq. 22 when the quadrature nodes come in complex conjugate pairs, i.e., $\mathfrak{i}\zeta_{j}=\mathfrak{i}s_{j}$ and $\mathfrak{i}\vartheta_{j}=-\mathfrak{i}s_{j}$ for all $j=1,\ldots,J$, and may overlap.
@@
3Evaluate the mappings at the quadrature nodes to obtain the data
$\left\{\Big{(}\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j}),\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j}),f(\mathfrak{i}\zeta_{j}),g(\mathfrak{i}\zeta_{j})\Big{)}\right\}_{j=1}^{J}\quad\text{and}\quad\left\{\Big{(}\boldsymbol{G}(\vartheta_{k}),f(\mathfrak{i}\vartheta_{k}),g(\mathfrak{i}\vartheta_{k})\Big{)}\right\}_{k=1}^{K},$
@@
4
5Compute the singular value decomposition
@@
6
7Compute the reduced-order model matrices according to
@@
[…]
@@
where $n$, $d$ and $h$ are defined as in eq. 20 for $f(s)=\alpha$ and $g(s)=\beta$; see [32, Def. 2]. If $\boldsymbol{C}_{\operatorname{p}}=\boldsymbol{0}$, and we take $\mu_{j}=\mathfrak{i}\zeta_{j}$ and $\lambda_{k}=\mathfrak{i}\vartheta_{k}$, then the matrices eq. 21, which appear in $\mathsf{soQuadpvBT}$, are related to the second-order Loewner and shifted Loewner matrices eq. 24 by diagonal scalings:
@@
In [32], the matrices eq. 24 are used to construct rational interpolants of second-order systems from data; see [32, Thm. 3] for further details. The input and output matrices of this rational interpolant are those in eq. 22c multiplied from the left and right, respectively, by the diagonal scalings given above. We also highlight that the derivation of the diagonally scaled Loewner matrices in the proof of Theorem 1 is different from that of [32, Thm. 3], which uses a parameterization of interpolating systems eq. 3 based on matrix equations. The tools and techniques used in the proof of Theorem 1 can be applied to derive this generalization of the Loewner framework from [32], as well. Significantly, this extends the applicability of the framework to second-order systems eq. 1 with the generalized proportional damping model eq. 5, as well as mixed position and velocity outputs.
Remark 4.
In [44], the authors re-derive the formulas from the dissertation [36] for the case of second-order velocity balanced truncation. In contrast to [44], our work not only considers the position-velocity balanced truncation approach but is also presented for a broader class of dynamical systems, namely multiple-input/multiple-output second-order systems with position as well as velocity outputs and with the generalized proportional damping model eq. 5. Additionally, we note that most of the following considerations in Sections 4 and 5 have not been discussed for the velocity balancing approach in [44].
@@ 4.1 Stability preservation
and the system matrices $\boldsymbol{M},\boldsymbol{D},\boldsymbol{K}$ are Hermitian positive definite ($HPD$). Obviously, a special case of eq. 25 is the system eq. 1 being state-space symmetric when the corresponding matrices are real valued. Interestingly, $\mathsf{soQuadpvBT}$ will preserve asymptotic stability if the data are generated by a Rayleigh-damped system that is Hermitian with $HPD$ mass, damping and stiffness matrices, and if the left and right quadrature nodes—those used to implicitly approximate $\boldsymbol{Q}_{\operatorname{v}}$ and $\boldsymbol{P}_{\operatorname{p}}$, respectively—are complex conjugate pairs of each other, i.e., $\mathfrak{i}\zeta_{i}=\mathfrak{i}s_{i}$ and $\mathfrak{i}\vartheta_{i}=-\mathfrak{i}s_{i}$. If the $s_{i}$’s are not all strictly positive or negative, then at least some of the nodes will be equivalent, i.e., $\zeta_{j}=s_{j}=-s_{k}=\vartheta_{k}$ for some $j,k$, which is not allowed under Theorem 1. To prove the stability result, we first derive formulae for the data matrices eq. 21 in this special case when the quadrature nodes underlying the approximations to $\boldsymbol{Q}_{\operatorname{v}}$ and $\boldsymbol{P}_{\operatorname{p}}$ come in complex conjugate pairs and may overlap, with $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0}$.
@@
Let the functions $d\colon\mathbb{C}\to\mathbb{C}$, $n\colon\mathbb{C}\to\mathbb{C}$ and $h\colon\mathbb{C}\to\mathbb{C}$ be defined as in eq. 20. Suppose the left and right quadrature rules in eqs. 16 and 17 are complex conjugate pairs, i.e., $\zeta_{i}=s_{i}$, $\vartheta_{i}=-s_{i}$ and $\eta_{i}=\varrho_{i}=\vartheta_{i}$ for $i=1,\ldots,N$ and $N=J=K$, and the quadrature rules are such that $d(\mathfrak{i}s_{i})\neq 0$ and $\boldsymbol{\varphi}(\mathfrak{i}s_{j})$ is nonsingular for all $i=1,\ldots,N$. Let the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ be given as in eqs. 16 and 17 and let the matrices $\mathbb{M},\mathbb{K}\in\mathbb{C}^{pK\times mJ}$, $\mathbb{B}_{\operatorname{u}}\in\mathbb{C}^{pK\times m}$ and $\mathbb{C}_{\operatorname{p}}\in\mathbb{C}^{p\times mJ}$ be as in eq. 21. Then, the matrices $\mathbb{M}$, $\mathbb{K}$, $\mathbb{B}_{\operatorname{u}}$ and $\mathbb{C}_{\operatorname{p}}$ are given block-entrywise by
$\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}$ =$=\begin{cases}\displaystyle-\frac{\eta_{k}\eta_{j}}{d(\mathfrak{i}s_{j})^{2}}\frac{\operatorname{d}}{\operatorname{d}s}\Big{(}d(s)\boldsymbol{G}(s)\Big{)}\bigg{|}_{s=\mathfrak{i}s_{j}}\left(\frac{\operatorname{d}}{\operatorname{d}s}h(s)\bigg{|}_{s=\mathfrak{i}s_{j}}\right)^{-1}&\text{if}\quad\mathfrak{i}s_{j}=-\mathfrak{i}s_{k},\\[14.22636pt] \displaystyle-\frac{\eta_{k}\eta_{j}}{d(-\mathfrak{i}s_{k})d(\mathfrak{i}s_{j})}\frac{d(-\mathfrak{i}s_{k})\boldsymbol{G}(-\mathfrak{i}s_{k})-d(\mathfrak{i}s_{j})\boldsymbol{G}(\mathfrak{i}s_{j})}{h(-\mathfrak{i}s_{k})-h(\mathfrak{i}s_{j})}&\text{if}\quad\mathfrak{i}s_{j}\neq-\mathfrak{i}s_{k},\end{cases}$ (26a)
$\mathbb{K}_{\boldsymbol{k},\boldsymbol{j}}$ =$=\begin{cases}\displaystyle\frac{\eta_{k}\eta_{j}}{d(\mathfrak{i}s_{j})^{2}}\frac{\operatorname{d}}{\operatorname{d}s}\Big{(}n(s)\boldsymbol{G}(s)\Big{)}\bigg{|}_{s=\mathfrak{i}s_{j}}\left(\frac{\operatorname{d}}{\operatorname{d}s}h(s)\bigg{|}_{s=\mathfrak{i}s_{j}}\right)^{-1}&\text{if}\quad\mathfrak{i}s_{j}=-\mathfrak{i}s_{k},\\[14.22636pt] \displaystyle\frac{\eta_{k}\eta_{j}}{d(-\mathfrak{i}s_{k})d(\mathfrak{i}s_{j})}\frac{n(-\mathfrak{i}s_{k})\boldsymbol{G}(-\mathfrak{i}s_{k})-n(\mathfrak{i}s_{j})\boldsymbol{G}(\mathfrak{i}s_{j})}{h(-\mathfrak{i}s_{k})-h(\mathfrak{i}s_{j})}&\text{if}\quad\mathfrak{i}s_{j}\neq-\mathfrak{i}s_{k},\end{cases}$ (26b)
$\big{(}\mathbb{B}_{\operatorname{u}}\big{)}_{\boldsymbol{k},:}$ $=\eta\boldsymbol{G}(-\mathfrak{i}s_{k})\quad\text{and}\quad\big{(}\mathbb{C}_{\operatorname{p}}\big{)}_{:,\boldsymbol{j}}=\eta\boldsymbol{G}(\mathfrak{i}s_{j}),$ (26c)
@@
$\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}=\boldsymbol{I}_{k,p}^{\mkern-1.5mu\mathsf{T}}\left({\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{M}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\right)\boldsymbol{I}_{j,m}=\eta_{k}\eta_{j}\boldsymbol{C}_{\operatorname{p}}\boldsymbol{\varphi}(\mathfrak{i}s_{j})^{-1}\boldsymbol{M}\boldsymbol{\varphi}(\mathfrak{i}s_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}$
@@
by the limit definition of the derivative. This proves eq. 26a when $\mathfrak{i}s_{j}=-\mathfrak{i}s_{k}$ by multiplication with $\eta_{k}\eta_{j}$. Using the same arguments, the other remaining formula eq. 26b can be shown for $\mathfrak{i}s_{j}=-\mathfrak{i}s_{k}$, which concludes the proof. ∎
@@
by the property eq. 25 of the underlying system. Moreover, $\mathbb{B}_{\operatorname{u}}=\mathbb{C}_{\operatorname{p}}^{\mathsf{H}}$ holds by construction. We claim that a reduced-order model eq. 3 computed via Algorithm 2 using these Hermitian data matrices has a similar symmetry structure as the original underlying system with $HPD$ mass, damping and stiffness matrices. Obviously ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}=\boldsymbol{I}_{r}$ is $HPD$ by construction. Because $\mathbb{M}$ is Hermitian, its singular value decomposition is so that the singular vectors satisfy ${\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}={\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}$. Then, for any $\boldsymbol{z}\in\mathbb{C}^{Nm}$, it holds that

### v2

@@ 3 Data-driven second-order balanced truncation
where $\boldsymbol{\phi}(s)\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\left(s\boldsymbol{M}+\boldsymbol{K}\right)$. To prove eqs. 22a and 22b, we use the resolvent identities eq. 19a and eq. 19b. Under the assumption that $h(\mathfrak{i}\vartheta_{k})\neq h(\mathfrak{i}\zeta_{j})$ for all $k$ and $j$, applying eq. 19a reveals that
@@
Replacing the intrusive quantities eq. 14 in Algorithm 1 with the data-based approximations eq. 21 results in a data-driven reformulation of $\mathsf{sopvBT}$ summarized in Algorithm 2, which we call second-order quadrature-based position-velocity balanced truncation ($\mathsf{soQuadpvBT}$). Outside of empirical knowledge used to design the damping coefficient functions $f$ and $g$, the method is entirely non-intrusive. As in [19], the quadrature-based approximations to the Gramians are never explicitly formed. Instead, they are only invoked implicitly to derive the data-based approximations eq. 21; the points in Algorithm 2 at which $\boldsymbol{G}_{\operatorname{p}}$ and $\boldsymbol{G}_{\operatorname{v}}$ are sampled are the quadrature nodes underlying these quadrature rules. While Algorithm 2 is formulated generally to include position and velocity outputs, as is the case with Theorem 1, both are not necessarily required; depending on the application of interest, only one of them is required. Lastly, Theorem 1 assumes that the sets of quadrature nodes used to implicitly approximate $\boldsymbol{P}_{\operatorname{p}}$ and $\boldsymbol{Q}_{\operatorname{v}}$ are distinct. In Theorem 2, we derive a special case of the formulae eq. 22 when the quadrature nodes come in complex conjugate pairs, i.e., $\mathfrak{i}\zeta_{j}=\mathfrak{i}s_{j}$ and $\mathfrak{i}\vartheta_{j}=-\mathfrak{i}s_{j}$ for all $j=1,\ldots,J$, and may overlap.
@@
2pt Output: Reduced-order system matrices ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}},{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}},{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ in eq. 3.
3
4Evaluate the mappings at the quadrature nodes to obtain the data
$\left\{\Big(\boldsymbol{G}_{\operatorname{p}}(\mathfrak{i}\zeta_{j}),\boldsymbol{G}_{\operatorname{v}}(\mathfrak{i}\zeta_{j}),f(\mathfrak{i}\zeta_{j}),g(\mathfrak{i}\zeta_{j})\Big)\right\}_{j=1}^{J}\quad\text{and}$
$\left\{\Big(\boldsymbol{G}(\vartheta_{k}),f(\mathfrak{i}\vartheta_{k}),g(\mathfrak{i}\vartheta_{k})\Big)\right\}_{k=1}^{K},$
@@
5
6Compute the singular value decomposition
@@
7
Compute the reduced-order model matrices according to
@@
${\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{}(s)$ $=f(s)\boldsymbol{I}_{r}+g(s){\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}$,
@@
where $n$, $d$ and $h$ are defined as in eq. 20 for $f(s)=\alpha$ and $g(s)=\beta$; see [29, Def. 2]. If $\boldsymbol{C}_{\operatorname{p}}=\boldsymbol{0}$, and we take $\mu_{j}=\mathfrak{i}\zeta_{j}$ and $\lambda_{k}=\mathfrak{i}\vartheta_{k}$, then the matrices eq. 21, which appear in $\mathsf{soQuadpvBT}$, are related to the second-order Loewner and shifted Loewner matrices eq. 24 by diagonal scalings:
@@
In [29], the matrices eq. 24 are used to construct rational interpolants of second-order systems from data; see [29, Thm. 3] for further details. The input and output matrices of this rational interpolant are those in eq. 22c multiplied from the left and right, respectively, by the diagonal scalings given above. We also highlight that the derivation of the diagonally scaled Loewner matrices in the proof of Theorem 1 is different from that of [29, Thm. 3], which uses a parameterization of interpolating systems eq. 3 based on matrix equations. The tools and techniques used in the proof of Theorem 1 can be applied to derive this generalization of the Loewner framework from [29], as well. Significantly, this extends the applicability of the framework to second-order systems eq. 1 with the generalized proportional damping model eq. 5, as well as mixed position and velocity outputs.
@@ 4.1 Stability preservation
and the system matrices $\boldsymbol{M},\boldsymbol{D},\boldsymbol{K}$ are Hermitian positive definite ($HPD$). Obviously, a special case of eq. 25 is the system eq. 1 being state-space symmetric when the corresponding matrices are real-valued. Interestingly, $\mathsf{soQuadpvBT}$ will preserve asymptotic stability if the data are generated by a Rayleigh-damped system that is Hermitian with $HPD$ mass, damping and stiffness matrices, and if the left and right quadrature nodes—those used to implicitly approximate $\boldsymbol{Q}_{\operatorname{v}}$ and $\boldsymbol{P}_{\operatorname{p}}$, respectively—are complex conjugate pairs of each other, i.e., $\mathfrak{i}\zeta_{i}=\mathfrak{i}s_{i}$ and $\mathfrak{i}\vartheta_{i}=-\mathfrak{i}s_{i}$. If the $s_{i}$’s are not all strictly positive or negative, then at least some of the nodes will be equivalent, i.e., $\zeta_{j}=s_{j}=-s_{k}=\vartheta_{k}$ for some $j,k$, which is not allowed under Theorem 1. To prove the stability result, we first derive formulae for the data matrices eq. 21 in this special case when the quadrature nodes underlying the approximations to $\boldsymbol{Q}_{\operatorname{v}}$ and $\boldsymbol{P}_{\operatorname{p}}$ come in complex conjugate pairs and may overlap, with $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0}$.
@@
Let the functions $d\colon\mathbb{C}\to\mathbb{C}$, $n\colon\mathbb{C}\to\mathbb{C}$ and $h\colon\mathbb{C}\to\mathbb{C}$ be defined as in eq. 20. Suppose the left and right quadrature rules in eqs. 16 and 17 are complex conjugate pairs, i.e., $\zeta_{i}=s_{i}$, $\vartheta_{i}=-s_{i}$ and $\xi_{i}=\varrho_{i}=\vartheta_{i}$ for $i=1,\ldots,N$ and $N=J=K$, and the quadrature rules are such that $d(\mathfrak{i}s_{i})\neq 0$ and $\boldsymbol{\varphi}(\mathfrak{i}s_{j})$ is nonsingular for all $i=1,\ldots,N$. Let the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ be given as in eqs. 16 and 17 and let the matrices $\mathbb{M},\mathbb{K}\in\mathbb{C}^{pK\times mJ}$, $\mathbb{B}_{\operatorname{u}}\in\mathbb{C}^{pK\times m}$ and $\mathbb{C}_{\operatorname{p}}\in\mathbb{C}^{p\times mJ}$ be as in eq. 21. Then, the matrices $\mathbb{M}$, $\mathbb{K}$, $\mathbb{B}_{\operatorname{u}}$ and $\mathbb{C}_{\operatorname{p}}$ are given block-entrywise by
$\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}$ $=\begin{cases}\displaystyle-\frac{\xi_{k}\xi_{j}}{d(\mathfrak{i}s_{j})^{2}}\frac{\operatorname{d}}{\operatorname{d}s}\Big(d(s)\boldsymbol{G}(s)\Big)\bigg|_{s=\mathfrak{i}s_{j}}\left(\frac{\operatorname{d}}{\operatorname{d}s}h(s)\bigg|_{s=\mathfrak{i}s_{j}}\right)^{-1}&\text{if}\quad\mathfrak{i}s_{j}=-\mathfrak{i}s_{k},\\[14.22636pt] \displaystyle-\frac{\xi_{k}\xi_{j}}{d(-\mathfrak{i}s_{k})d(\mathfrak{i}s_{j})}\frac{d(-\mathfrak{i}s_{k})\boldsymbol{G}(-\mathfrak{i}s_{k})-d(\mathfrak{i}s_{j})\boldsymbol{G}(\mathfrak{i}s_{j})}{h(-\mathfrak{i}s_{k})-h(\mathfrak{i}s_{j})}&\text{if}\quad\mathfrak{i}s_{j}\neq-\mathfrak{i}s_{k},\end{cases}$ (26a)
$\mathbb{K}_{\boldsymbol{k},\boldsymbol{j}}$ $=\begin{cases}\displaystyle\frac{\xi_{k}\xi_{j}}{d(\mathfrak{i}s_{j})^{2}}\frac{\operatorname{d}}{\operatorname{d}s}\Big(n(s)\boldsymbol{G}(s)\Big)\bigg|_{s=\mathfrak{i}s_{j}}\left(\frac{\operatorname{d}}{\operatorname{d}s}h(s)\bigg|_{s=\mathfrak{i}s_{j}}\right)^{-1}&\text{if}\quad\mathfrak{i}s_{j}=-\mathfrak{i}s_{k},\\[14.22636pt] \displaystyle\frac{\xi_{k}\xi_{j}}{d(-\mathfrak{i}s_{k})d(\mathfrak{i}s_{j})}\frac{n(-\mathfrak{i}s_{k})\boldsymbol{G}(-\mathfrak{i}s_{k})-n(\mathfrak{i}s_{j})\boldsymbol{G}(\mathfrak{i}s_{j})}{h(-\mathfrak{i}s_{k})-h(\mathfrak{i}s_{j})}&\text{if}\quad\mathfrak{i}s_{j}\neq-\mathfrak{i}s_{k},\end{cases}$ (26b)
$\big(\mathbb{B}_{\operatorname{u}}$ $\big)_{\boldsymbol{k},:}=\xi_{k}\boldsymbol{G}(-\mathfrak{i}s_{k})\quad\text{and}\quad\big(\mathbb{C}_{\operatorname{p}}\big)_{:,\boldsymbol{j}}=\xi_{j}\boldsymbol{G}(\mathfrak{i}s_{j}),$ (26c)
@@
$\mathbb{M}_{\boldsymbol{k},\boldsymbol{j}}=\boldsymbol{I}_{k,p}^{\mkern-1.5mu\mathsf{T}}\left({\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{M}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\right)\boldsymbol{I}_{j,m}=\xi_{k}\xi_{j}\boldsymbol{C}_{\operatorname{p}}\boldsymbol{\varphi}(\mathfrak{i}s_{j})^{-1}\boldsymbol{M}\boldsymbol{\varphi}(\mathfrak{i}s_{j})^{-1}\boldsymbol{B}_{\operatorname{u}}$
@@
by the limit definition of the derivative. This proves eq. 26a when $\mathfrak{i}s_{j}=-\mathfrak{i}s_{k}$ by multiplication with $\xi_{k}\xi_{j}$. Using the same arguments, the other remaining formula eq. 26b can be shown for $\mathfrak{i}s_{j}=-\mathfrak{i}s_{k}$, which concludes the proof. ∎
@@
by the property eq. 25 of the underlying system. Moreover, $\mathbb{B}_{\operatorname{u}}=\mathbb{C}_{\operatorname{p}}^{\mathsf{H}}$ holds by construction. We claim that a reduced-order model eq. 3 computed via Algorithm 2 using these Hermitian data matrices has a similar symmetry structure as the original underlying system with $HPD$ mass, damping and stiffness matrices. Obviously ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{}=\boldsymbol{I}_{r}$ is $HPD$ by construction. Because $\mathbb{M}$ is Hermitian, its singular value decomposition is such that the singular vectors satisfy ${\check{\boldsymbol{U}\mkern 2.0mu}\mkern-2.0mu}{}={\check{\boldsymbol{Y}\mkern 2.0mu}\mkern-2.0mu}{}$. Then, for any $\boldsymbol{z}\in\mathbb{C}^{Nm}$, it holds that